OpenStudy (anonymous):

The distance between point (4,0) and point B is 7. The coordinates of point B could be any of the following except?

OpenStudy (anonymous):

(4,7) (-3,0) (11,0) (3,4) (4,-7)

OpenStudy (anonymous):

11,0?

OpenStudy (kamille):

correct

TheSmartOne (thesmartone):

\(\Large\color{red}{ Distance=\sqrt {(x_2-x_1)^2+(y_2-y_1)^2}}\)

OpenStudy (anonymous):

okay so wouldn't it be 4,-7

TheSmartOne (thesmartone):

\(\large 7=\sqrt {(11-4)^2+(0-0)^2}\) \(\large 7=\sqrt {(7)^2+(0)^2}\) \(\large 7=7\)

TheSmartOne (thesmartone):

So yes your asnwer would be (11,0)

OpenStudy (anonymous):

I know the question ask which one it wouldn't be

TheSmartOne (thesmartone):

Oh in that case (11,0) is not your answer...

OpenStudy (anonymous):

lol yeah I'm trying it right now

TheSmartOne (thesmartone):

(4,7) is not teh answer either www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=4&val2=0&val3=4&val4=7&rb1=dist

TheSmartOne (thesmartone):

(-3,0) is not it either www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=4&val2=0&val3=-3&val4=0&rb1=dist

TheSmartOne (thesmartone):

(3,4) is your asnwer www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=4&val2=0&val3=3&val4=4&rb1=dist

TheSmartOne (thesmartone):

nad neiter is (4,-7) www.mathportal.org/calculators/analytic-geometry/distance-and-midpoint-calculator.php?val1=4&val2=0&val3=4&val4=-7&rb1=dist

TheSmartOne (thesmartone):

\(\Huge\color{blue}{(3,4)~is~your~answer!!}\)

OpenStudy (anonymous):

okay thank you so much!!! merci beaucoup!!

TheSmartOne (thesmartone):

np